what are the side lengths of a rectangle whos perimeter is 66 and its area is 216

Answer by xdragonfight(116)

2L+2W=66
LxW=216

so start with 2L+2W=66 lets solve for W
2L+2W=66
-2L -2L
2W=66-2L
divide everything by 2
2w/2=w 66/2=33 -2L/2=-L
so new equation is w=33-L ok lets start on the 2nd equation
LxW=216
replace W with (33-L)
Lx(33-L)=216
-L^2+33L=216
so L=9
ok lets find W now
2(9)+2W=66
2W=48
W=24
L=9
24x9=216